OpenStudy (anonymous):

change 7 5/14 to a decimal & round to the answer two decimal places

OpenStudy (anonymous):

\[7\frac{ 5 }{ 14 }\]

OpenStudy (anonymous):

7.35

OpenStudy (kirbykirby):

Never mind i realized there's no multiplication sign..

OpenStudy (kirbykirby):

SO just do 14*7 = 98 98 + 5 = 103 Then do 103/14

OpenStudy (anonymous):

13/200?

OpenStudy (kirbykirby):

And to round your answer to 2 decimal places I would put 7.36 not 7.35 because your number is 7.35714... and the number after the "5" is a "7" so you should round up